According to a report by Scarborough Research, the monthly average cell phone bill is $73.55. Suppose local monthly household cell phone bills are normally distributed with a standard deviation of $11.30. Between what monthly costs do 95% of households fall according to this report?
Answer by ikleyn(52781)

.
According to a report by Scarborough Research, the monthly average cell phone bill is $73.55.
Suppose local monthly household cell phone bills are normally distributed with a standard
deviation of $11.30. Between what monthly costs do 95% of households fall according to this report?
~~~~~~~~~~~~~~~~~~~~~~~~~

The "empirical rule" in a normal distribution states that approximately 95% data points 
fall within two standard deviations.


So, in this problem the lower bound is m - 2d = 73.55 - 2*11.30 = 50.95 dollars;

                    the upper bound is m + 2d = 73.55 + 2*11.30 = 96.15 dollars.

Solved.